OpenStudy (anonymous):

What is the focal width of (-1/12)x^2=y?

OpenStudy (jdoe0001):

what's a focal width?

OpenStudy (anonymous):

length of a vertical or horizontal line that passes through the focus and touches the parabola on each end

OpenStudy (anonymous):

it says |4a|

OpenStudy (jdoe0001):

4a?

OpenStudy (anonymous):

I don't know what it means that's why I asked lol

OpenStudy (jdoe0001):

heh, I mean, I know what you're saying, I was looking for a good graphic :)

OpenStudy (anonymous):

nevermind the 4a that was an example i thought it was like an equation for focal width

OpenStudy (jdoe0001):

but is not 4a as far as I know

OpenStudy (jdoe0001):

one sec

OpenStudy (anonymous):

okay! :)

OpenStudy (anonymous):

i think it's either 48 or 12

OpenStudy (anonymous):

12

OpenStudy (jdoe0001):

the "c" or so-called focal point is \(c =\cfrac{1}{4a} \)

OpenStudy (jdoe0001):

like the (h,k+a) in the picture http://www.boredofstudies.org/wiki/images/6/6f/Parabola.png

OpenStudy (jdoe0001):

so, from \(y = -\cfrac{1}{12} x^2\) the "a" component would be coefficient of "x"

OpenStudy (anonymous):

so is 12 a?

OpenStudy (anonymous):

or is a 3?

OpenStudy (anonymous):

I'm confused lol

OpenStudy (jdoe0001):

well, you have a negative \(-\frac{1}{12}\) that'd give you a -3

OpenStudy (anonymous):

oh true but if c=1/4a then wouldn't c=1/12??

OpenStudy (anonymous):

with a negative lol

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

but the only possible choices are 12 and 48